# Parallel Test-Time Scaling for Latent Reasoning Models

### v1

@@ 6.2 Analysis on Diversity
To gain deeper insight into how different sampling methods influence the reasoning process over time, we further analyze diversity at each individual reasoning step. As illustrated in Figure 4, we find that MC-dropout maintains relatively similar diversity across steps under same $p$, reflecting its adaptive nature; its stochasticity is inherently tied to the model’s own assessment, allowing for consistent and "self-aware" exploration; whereas AGN exhibits pronounced fluctuations, due to its use of a fixed scale $\sigma$ that perturbs latent vectors of varying magnitudes unevenly across steps, ultimately resulting in inconsistent stochastic influence. These observations align with the theoretical mechanisms of each method: MC-dropout modulates its stochastic influence based on the model’s epistemic uncertainty, thereby supporting more effective exploration; whereas AGN, while allowing diversity to be increased by manually setting $\sigma$, introduces noise in a less adaptive manner, leading to variable impact across reasoning steps.
@@
Figure 6: Test-time scaling results with majority voting, best-of-NN, and beam search for latent reasoning. Accuracy(%) versus compute budget NN under three aggregation strategies on (6(a)) GSM8K-Test, (6(b)) GSM8K-Hard, and (6(c)) MultiArith. Results are reported as the mean over three runs.
@@ 7.2 Ablation Studies
To understand the contribution of each component in LatentRM, we conduct ablation studies under the Best-of-$N$ setting with $N=8$ on GSM-Test and GSM-Hard (Table 1). Each variant disables one design element of LatentRM while keeping all other configurations identical. (1) Removing the step-wise contrastive loss (w/o contrastive) causes a noticeable drop, showing that relative supervision among concurrent thoughts provides stronger learning signals than isolated binary labels. (2) Excluding stochastic rollouts (w/o stochastic rollouts), where each thought is labeled only by the final trajectory correctness instead of intermediate estimates, leads to a further decline, highlighting the importance of Monte Carlo estimation for reliable thought-level annotation. (3) Using an untrained random scalar head (random head) performs even worse—below Majority Voting—indicating that the gain stems from learned evaluation rather than architectural modification. Together, these findings verify that LatentRM effectively learns to assess latent trajectories and is crucial for successful aggregation in latent test-time scaling.
@@ 8 Conclusion and Future Work
Building on this foundation, future work will seek to integrate sampling and aggregation into a reinforcement learning framework, where latent trajectories are optimized through iterative feedback and reward shaping. Such a formulation could transform test-time scaling from a static inference procedure into an adaptive reasoning process—capable of learning exploration strategies, adjusting compute allocation dynamically, and generalizing across tasks.
Limitation
This work provides an effective framework for parallel test-time scaling in latent reasoning models, but there are certain limitations: (1) Engineering considerations for real-time deployment. The framework shows consistent performance and generality across benchmarks and compute budgets, but real-time use may require additional engineering optimization. Nevertheless, as the core contribution is a proof-of-concept and in-depth analysis of parallel scaling for latent reasoning models, these engineering considerations do not impact the validity or the key insights presented in this study. (2) Hyperparameter sensitivity. Our sampling strategies demonstrate strong scaling results with controllable randomness. But optimal performance depends on tuning the dropout rate ($p$) and noise scale ($\sigma$). While this could be seen as an additional configuration burden, we note that the hyperparameter search process is relatively straightforward and does not require extra model training or complex optimization. The main contribution of this work remains focused on the theoretical framework and empirical performance, thus the potential overhead from hyperparameter tuning is minor compared to the overall benefits.
Ethics Statement
All data, models and other related experimental artifacts used in this study are publicly released and accessible under permissive licenses for research purposes. The research procedures, data handling, and analyses do not involve harm to individuals, communities, or the environment, and no personal or sensitive information is included.
Appendix A Theoretical Perspectives on Sampling in Latent Space
@@
can be interpreted as an approximate Bayesian method, where stochastic forward passes approximate the posterior predictive distribution given the training dataset $D$ Gal and Ghahramani (2016). Let the predictive distribution be
@@
Appendix B Derivation of Cumulative Scoring
@@
Appendix C Inference Procedures for Aggregation
@@
Appendix D Training Configuration for LatentRM
We train a Latent Reward Model (LatentRM) for COCONUT using the step-wise contrastive learning objective introduced in Section 4.2. model initialized from COOCNUT checkpoint, a GPT-2 backbone with 124 million parameters Radford et al. , with architecture modifications detailed therein.
@@ Appendix D Training Configuration for LatentRM
The training data consists of samples generated from the GSM8K training set (385K examples). For each input problem, we sample $N=8$ reasoning trajectories via MC-dropout with dropout probability $p=0.2$ from COCONUT, where $p$ was tuned for optimal performance as mentioned in Section 6.1.

### v3

@@ 6.2 Analysis on Diversity
To examine how sampling affects reasoning over time, we analyze diversity at each step. As shown in Figure 3, MC-dropout maintains similar diversity across steps for a given $p$, reflecting its adaptivity: stochasticity follows the model’s own uncertainty, enabling consistent exploration. In contrast, AGN shows pronounced fluctuations because a fixed $\sigma$ perturbs latent vectors of varying scales unevenly, yielding inconsistent stochastic influence. These patterns match each method’s mechanism: MC-dropout modulates noise by epistemic uncertainty, supporting effective exploration, whereas AGN, though diversity can be increased via $\sigma$, injects less adaptive noise with variable impact across steps.
@@
Figure 5: Test-time scaling results with majority voting, best-of-NN, and beam search for latent reasoning. Accuracy(%) versus compute budget NN under three aggregation strategies on (5(a)) GSM8K-Test, (5(b)) GSM8K-Hard, and (5(c)) MultiArith. Results are reported as the mean over three runs.
@@ 7.2 Ablation Studies
To understand the contribution of each component in LatentRM, we conduct ablation studies under the Best-of-$N$ setting with $N=8$ on GSM-Test and GSM-Hard (). Each variant disables one design element of LatentRM while keeping all other configurations identical. (1) Removing the step-wise contrastive loss (w/o contrastive) causes a noticeable drop, showing that relative supervision among concurrent thoughts provides stronger learning signals than isolated binary labels. (2) Excluding stochastic rollouts (w/o stochastic rollouts), where each thought is labeled only by the final trajectory correctness instead of intermediate estimates, leads to a further decline, highlighting the importance of Monte Carlo estimation for reliable thought-level annotation. (3) Using an untrained random scalar head (random head) performs even worse—below Majority Voting—indicating that the gain stems from learned evaluation rather than architectural modification. Together, these findings verify that LatentRM effectively learns to assess latent trajectories and is crucial for successful aggregation in latent test-time scaling.
7.3 More Analysis
Due to space constraints, further extended analyses on Comparison of Latent and Explicit Reasoning with TTS and Analysis on LatentRM under Variable-Length Setting are provided in Appendix F.1 and Appendix F.2.
@@ 8 Conclusion and Future Work
Building on this foundation, future research should explore two key directions: (1) integrating sampling and aggregation into a reinforcement learning framework to optimize latent trajectories through iterative feedback and reward shaping; and
Appendix A Soft and Latent Reasoning
Soft Thinking [llms_single_threaded_2025, SoftThinking] and Latent Reasoning [COCONUT, COLAR] represent two distinct paradigms for reasoning in large language models. Soft Thinking operates in the token probability space, where "soft tokens" are created by mixing token embeddings according to probability distributions over the vocabulary. This approach effectively retains the vocabulary structure while enabling continuous interpolation between discrete tokens [SoftThinking]. In contrast, Latent Reasoning works directly with pure latent vectors (hidden states) that are not constrained to be mixtures of token embeddings, operating in a more abstract and unconstrained continuous space [COCONUT].
Wu et al. [wu_parallel_2025] reveal that Soft Thinking suffers from a greedy pitfall: LLMs predominantly rely on the highest-probability token, suppressing alternative reasoning paths. Their Stochastic Soft Thinking addresses this by introducing controlled randomness (via Gumbel-Softmax) in soft embedding composition. In contrast, our Parallel TTS for Latent Reasoning operates in pure latent vector space, unconstrained by token embeddings or vocabulary structure. We combine latent sampling (MC-Dropout or AGN) with a Latent Reward Model to enable parallel exploration and aggregation, discovering reasoning patterns unbound by vocabulary structure.
Appendix B Theoretical Perspectives on Sampling in Latent Space
@@
can be interpreted as an approximate Bayesian method, where stochastic forward passes approximate the posterior predictive distribution given the training dataset $D$ [gal_dropout_2016]. Let the predictive distribution be
@@
Appendix C Derivation of Cumulative Scoring
@@
Appendix D Inference Procedures for Aggregation
@@
Appendix E Additional Results on Harder Benchmarks
We further evaluate our method on harder benchmarks, GPQA and AIME. As shown in Figure 6, our method induces consistent performance gains even on these harder benchmarks, which demonstrates the general effectiveness of the proposed approach. However, absolute performance remains limited. We attribute to two factors: (a) small model size (e.g., GPT-2-124M, Llama-3.2-1B), and (b) the inherent challenges of the "latent reasoning" framework, which is still an evolving paradigm and struggle on advanced mathematics (AIME) and Ph.D.-level benchmark (GPQA).
(a) GPQA
(b) AIME
Figure 6:
Appendix F More Analysis
F.1 Comparison of Latent and Explicit Reasoning with TTS
To investigate the synergy between latent reasoning and Test-Time Scaling (TTS), we compare No-CoT (direct answering), explicit CoT, and COCONUT (latent reasoning), evaluating each with and without parallel TTS ($N=4$). We measure computational overhead by the number of decoding steps relative to the No-CoT baseline.
Method GSM-Test GSM-Hard Compute
NO-COT
deterministic 16.5 4.3 $\times$1
Coverage@4 22.3 5.5 $\times$4
COT
deterministic 42.9 9.8 $\times$30
Coverage@4 50.7 11.6 $\times$120
COCONUT
deterministic 34.0 7.0 $\times$6
Coverage@4 46.7 9.5 $\times$24
Best-of-4 34.5 7.5 $\times$24
Average compute across Test and Hard subsets.
Table 2:
As summarized in Table 2, latent reasoning combined with TTS demonstrates superior efficiency, achieving competitive performance with significantly lower computational costs than explicit CoT. Specifically, COCONUT with TTS outperforms the explicit CoT baseline on the same backbone while requiring fewer decoding steps. Furthermore, while a performance gap persists between the Best-of-$N$ with LatentRM and the Coverage@$N$ upper bound, integrating LatentRM provides consistent gains. These results suggest that latent reasoning, augmented by TTS, preserves the efficiency advantages of continuous-space processing while matching or exceeding the accuracy of discrete reasoning.
F.2 Analysis on LatentRM under Variable-Length Setting
Method Accuracy (%)
Deterministic 32.44
Best-of-4 33.81
Best-of-8 33.95
Best-of-16 34.04
Table 3:
Latent reasoning models may generate trajectories of varying lengths, we thus conduct an evaluation of LatentRM in a variable-length setting. We modify the COCONUT paradigm by introducing stochasticity into the trajectory generation process, sampling the number of reasoning steps uniformly from 4 to 6, instead of fixing the trajectory length to 6. To mitigate potential length bias in trajectory selection, we utilize the mean step-wise reward, defined as the total reward divided by the step count, to normalize scores. Experimental results in Table 3 indicate that best-of-$N$ reranking based on average reward consistently outperforms the deterministic baseline, suggesting that LatentRM can effectively handle variable-length latent reasoning.
Appendix G Training Configuration for LatentRM
We train a Latent Reward Model (LatentRM) for COCONUT using the step-wise contrastive learning objective introduced in Section 4.2. model initialized from COOCNUT checkpoint, a GPT-2 backbone with 124 million parameters [radford_language_nodate], with architecture modifications detailed therein.
@@ Appendix G Training Configuration for LatentRM
The training data consists of samples generated based on GSM8K training set (385K examples). For each input problem, we sample $N=8$ reasoning trajectories via MC-dropout with dropout probability $p=0.2$ from COCONUT, where $p$ was tuned for optimal performance as mentioned in Section 6.1.
@@
Appendix H Algorithmic Specification of Latent Sampling Procedures
This section presents the pseudocode for the stochastic sampling procedures used in latent test-time scaling. We detail the inference-time execution of MC-Dropout and AGN for generating multiple latent reasoning trajectories.
Algorithm 1 Sampling with MC-Dropout
1:Input sequence $x$; model $f_{\theta}^{\text{latent}}$; dropout rate $p$; sample count NN; stop-thinking condition STOP()($\cdot$).
2:Prefill: Encode $x$ and initialize kv-cache.
3:for $n=1$ to NN do
4:  $t\leftarrow 0$
5:  while not $\text{STOP}(h^{(n)}_{1:t})$ do
6:    $m_{t}^{(n)}\sim\text{Bernoulli}(1-p)$ $\triangleright$ Sample dropout mask
7:    $\theta_{t}^{(n)}\leftarrow\theta\odot m_{t}^{(n)}$ $\triangleright$ Masked Subnetwork
8:    $h^{(n)}_{t+1}\leftarrow f_{\bm{\theta}^{(n)}}(x,h^{(n)}_{1:t}) \triangleright$ Forward pass
9:    $t\leftarrow t+1$
10:  end while
11:  Store trajectory $\mathbf{h}^{(n)}\leftarrow[h^{(n)}_{1},\ldots,h^{(n)}_{t}$]
12:  Generate final answers $y^{(n)}\leftarrow f_{\theta}(x,\mathbf{h}^{(n)})$
13:end for
Algorithm 2 Sampling with AGN
1:Input sequence $x$; model $f_{\theta}^{\text{latent}}$; noise scale $\sigma$; sample count NN; stop-thinking condition STOP()($\cdot$).
2:Prefill: Encode $x$ and initialize kv-cache.
3:for $n=1$ to NN do
4:  $t\leftarrow 0$
5:  Initialize latent state $h^{(n)}_{1:0}\leftarrow\emptyset$
6:  while not $\text{STOP}(h^{(n)}_{1:t})$ do
7:    $\epsilon_{t}^{(n)}\sim\mathcal{N}(0,\sigma^{2}I)$ $\triangleright$ Sample Gaussian noise
8:    $\tilde{h}_{t}^{(n)}\leftarrow h_{t}^{(n)}+\epsilon_{t}^{(n)}$ $\triangleright$ Perturb latent thought
9:    $h^{(n)}_{t+1}\leftarrow f_{\theta}^{\text{latent}}(x,\tilde{h}^{(n)}_{1:t}) \triangleright$ Forward pass
10:    $t\leftarrow t+1$
11:  end while
12:  Store trajectory $\mathbf{h}^{(n)}\leftarrow[h^{(n)}_{1},\ldots,h^{(n)}_{t}$]
13:  Generate final answers $y^{(n)}\leftarrow f_{\theta}(x,\mathbf{h}^{(n)})$
14:end for